# ProtTeX: Structure-In-Context Reasoning and Editing of Proteins with Large Language Models

### v2

@@ 3.2.   ProtTeX Enables Structure-Involved Reasoning for Proteins
LLMs have demonstrated remarkable reasoning capabilities, particularly through Chain-of-Thought (CoT) reasoning, which involves decomposing complex problems into sequential and logical steps. This method enables models to generate coherent and contextually relevant responses. CoT reasoning has been successfully implemented in several prominent LLMs, including OpenAI-o1 58 and DeepSeek-R1 53. In multimodal settings, Chain-of-Thought reasoning allows models to effectively integrate and analyze diverse data types—such as text, images, and audio—by establishing meaningful connections among them. Although recent studies 59, 60 have begun exploring the integration of CoT reasoning into autoregressive image generation, the potential applications of such reasoning techniques in the biological sciences remain largely unexplored.
@@
Given the demonstrated effectiveness of Chain-of-Thought (CoT) reasoning in enhancing functional prediction tasks, we further investigate whether an alternative CoT strategy—beginning with sequence analysis followed by structure prediction—can similarly improve performance in structure prediction tasks. Traditional protein structure prediction inherently involves significant computational challenges, typically requiring specialized SE(3)-invariant architectures and resource-intensive, multi-phase training paradigms. In contrast, our model accomplishes this task solely through next-token prediction. Since LLMs are decoder-only generative models, we explored whether the model itself can intrinsically assess the quality of the generated protein structures. To this end, we employed perplexity (PPL)—a widely used metric in natural language processing (NLP)—which measures how effectively a probabilistic model predicts given samples 61. Empirical validation conducted on 500 randomly selected proteins from the PSPD test set revealed a statistically significant negative correlation between TM-score and output perplexity, as illustrated in Figure 4B. This correlation suggests that minimizing perplexity could enhance prediction accuracy. Therefore, we propose a simple sampling strategy termed "Beam Search with Lowest PPL." Detailed descriptions of this method are provided in Section 2.3.3. As demonstrated in Figure 4C and Figure 4D, our proposed sampling strategy improves structural prediction accuracy, highlighting the effectiveness of sampling-based approaches in enhancing predictive performance. These findings open promising directions for future research aimed at developing more efficient and precise sampling strategies, potentially advancing the application of autoregressive models in multimodal scenarios.
@@ 3.3.   Transaction of Language Decoding Techniques for Protein Structure Sampling
In LLMs, sampling is a critical step during text generation, with temperature parameters often controlling the balance between accuracy and diversity 62. Similarly, this principle may be beneficial for generating protein structures, where accurately modeling conformational variability is essential. Protein conformational diversity allows proteins to adopt distinct structural states under varying physiological conditions, significantly impacting protein behavior, ligand binding, and allosteric regulation 63. To explore this concept, we conducted zero-shot experiments to evaluate our model’s capability for generating multiple protein conformations.
We selected a set of nine proteins known to exhibit conformational variability. Specifically, three proteins (KaiB, Mad2, and RfaH) were previously studied by AlphaFold using a multiple sequence alignment (MSA) clustering method 64. The remaining six proteins, MinE, EhCaBP, DDX19, IMPase, Thioesterase, and Capsid Protein, which exhibit co-evolved residue pairs, are classified into Category 1 by W. Schafer et al. 48. Since our model was not explicitly trained for this task, we defined successful sampling as the ability to sample two distinct conformations, each with a TM-score above 0.7.
We utilized the nucleus sampling strategy 45, a widely recognized and effective approach for text generation. Specifically, we set the temperature parameter to 0.7 and the top-p value to 0.4, ensuring controlled generation diversity and quality. For each protein, we generated 100 samples and selected the pairs that exhibitd the highest structural similarity to the two target conformations. Our model successfully sampled 6/9 proteins. As shown in Figure 5, all three proteins identified by AFcluster were successfully sampled by our model, although their secondary structures remained suboptimal. The other three successful cases are presented in the Supplementary Figure S1.
@@ 4.   Concluding Remarks
Our current model may exhibit slight performance gaps in certain tasks compared to task-specific or domain-expert models, such as ESMFold2, 1. Notably, ESMFold’s success is largely attributed to its massive parameter count and extensive training dataset, which surpass even those of Llama3 8B in scale. Considering the well-known scaling laws of LLMs, we can anticipate continuous performance improvements across various tasks with the development of larger LLMs and more biological data. Beyond scaling parameters, well-established paradigms in the LLM domain, such as reinforcement learning-based alignment65 and inference-time self-improvement66, can be systematically applied to enhance ProtTeX’s performance across various protein-related tasks. Exploring these approaches will be a key focus for our future research.
@@ A.2.   Dataset
UniProt42 is a comprehensive and widely used database for protein annotation, enriched with detailed functional information and corresponding protein sequences. Many protein-related question-answering datasets focus on protein functionality. For example, Mol-Instruction31 and ProteinLMBench 20 are both constructed based on the extensive information provided by UniProt.To establish a robust protein structure database, we first construct a dataset based on UniProt. Specifically, we collect the clustered AlphaFold Protein Structure Database (AFDB) v4 dataset39, which includes 2.27 million single-chain structures predicted before July 25, 2022. Additionally, we extract protein structures from the Swiss-Prot database, filtering and curating the dataset released in May 2022, resulting in 541,327 single-chain structures. Furthermore, we incorporate experimentally determined structures from RCSB PDB41, including 551,957 single-chain structures released before October 13, 2021. In total, our dataset comprises 3.36 million protein sequences paired with structural information, forming a comprehensive resource for protein structure analysis and model training.
@@
Protein Structure Analysis Dataset(PSAD). We curate the dataset from proteinLMBench 20. The original questions required the model to generate descriptions of protein subunit composition, including the organization and interactions between subunits based on the given protein sequence. To enhance the model’s capabilities, we redesigned the prompt to enable protein structure generation following the completion of structural analysis, as illustrated in Figure 1C. This modification allows the model to achieve multimodal Chain-of-Thought (CoT) functionality.
@@ A.3.   Training Details
In our experiments, we employ both full-parameter fine-tuning and Low-Rank Adaptation (LoRA) 49 to optimize the performance of LLMs. To determine the most effective approach, we conduct ablation studies, which reveal that LoRA significantly underperforms compared to full-parameter fine-tuning in terms of task-specific metrics. Consequently, we adopt full-parameter fine-tuning as our primary training strategy. For optimization, we use the AdamW optimizer with a weight decay of 0.1. During training, we apply a cosine annealing learning rate scheduler, gradually decreasing the learning rate from 5e-6 to 1e-7, complemented by a warm-up phase with a ratio of 0.01. Training is conducted on the entire dataset with a batch size of 3 per device over 4 epochs, requiring approximately 5 days to complete on a cluster of 16 NVIDIA A100 GPUs.
@@ A.4.   Ablation details
Ablation on training strategy. We compare the LoRA (Low-Rank Adaptation) 49 training method with full-parameter fine-tuning. For full-parameter fine-tuning, we adopt the configuration previously described in Section A.3. For LoRA implementation, we configure the following parameters: LoRA target set to ‘all’, LoRA rank = 8, and alpha = 16. The learning rate is set higher than that of full-parameter fine-tuning, initialized at 1e-4 and gradually decayed to 1e-6, while all other hyperparameters remain consistent with the full-parameter fine-tuning setup. Our experimental results indicate that LoRA training exhibited significantly slower convergence in validation loss compared to full-parameter fine-tuning, as illustrated in the supplementary figures S2. Moreover, with the application of DeepSpeed Zero Stage 3 technology, LoRA does not demonstrate significant advantages in terms of memory usage or training speed across multiple GPUs compared to full-parameter fine-tuning. Consequently, we adopt full-parameter fine-tuning as our standard training paradigm in subsequent experiments.
@@
Figure S1: Multi-conformation sampling for fold-switching proteins in W. Schafer et al. 48.

### v3

@@ 3.2.   ProtTeX Enables Structure-Involved Reasoning for Proteins
LLMs have demonstrated remarkable reasoning capabilities, particularly through Chain-of-Thought (CoT) reasoning, which involves decomposing complex problems into sequential and logical steps. This method enables models to generate coherent and contextually relevant responses. CoT reasoning has been successfully implemented in several prominent LLMs, including OpenAI-o1 57 and DeepSeek-R1 52. In multimodal settings, Chain-of-Thought reasoning allows models to effectively integrate and analyze diverse data types—such as text, images, and audio—by establishing meaningful connections among them. Although recent studies 58, 59 have begun exploring the integration of CoT reasoning into autoregressive image generation, the potential applications of such reasoning techniques in the biological sciences remain largely unexplored.
@@
Given the demonstrated effectiveness of Chain-of-Thought (CoT) reasoning in enhancing functional prediction tasks, we further investigate whether an alternative CoT strategy—beginning with sequence analysis followed by structure prediction—can similarly improve performance in structure prediction tasks. Traditional protein structure prediction inherently involves significant computational challenges, typically requiring specialized SE(3)-invariant architectures and resource-intensive, multi-phase training paradigms. In contrast, our model accomplishes this task solely through next-token prediction. Since LLMs are decoder-only generative models, we explored whether the model itself can intrinsically assess the quality of the generated protein structures. To this end, we employed perplexity (PPL)—a widely used metric in natural language processing (NLP)—which measures how effectively a probabilistic model predicts given samples 60. Empirical validation conducted on 500 randomly selected proteins from the PSPD test set revealed a statistically significant negative correlation between TM-score and output perplexity, as illustrated in Figure 4B. This correlation suggests that minimizing perplexity could enhance prediction accuracy. Therefore, we propose a simple sampling strategy termed "Beam Search with Lowest PPL." Detailed descriptions of this method are provided in Section 2.3.3. As demonstrated in Figure 4C and Figure 4D, our proposed sampling strategy improves structural prediction accuracy, highlighting the effectiveness of sampling-based approaches in enhancing predictive performance. These findings open promising directions for future research aimed at developing more efficient and precise sampling strategies, potentially advancing the application of autoregressive models in multimodal scenarios.
@@ 3.3.   Transaction of Language Decoding Techniques for Protein Structure Sampling
In LLMs, sampling is a critical step during text generation, with temperature parameters often controlling the balance between accuracy and diversity 61. Similarly, this principle may be beneficial for generating protein structures, where accurately modeling conformational variability is essential. Protein conformational diversity allows proteins to adopt distinct structural states under varying physiological conditions, significantly impacting protein behavior, ligand binding, and allosteric regulation 62. To explore this concept, we conducted zero-shot experiments to evaluate our model’s capability for generating multiple protein conformations.
We selected a set of nine proteins known to exhibit conformational variability. Specifically, three proteins (KaiB, Mad2, and RfaH) were previously studied by AlphaFold using a multiple sequence alignment (MSA) clustering method 63. The remaining six proteins, MinE, EhCaBP, DDX19, IMPase, Thioesterase, and Capsid Protein, which exhibit co-evolved residue pairs, are classified into Category 1 by W. Schafer et al. 25. Since our model was not explicitly trained for this task, we defined successful sampling as the ability to sample two distinct conformations, each with a TM-score above 0.7.
We utilized the nucleus sampling strategy 46, a widely recognized and effective approach for text generation. Specifically, we set the temperature parameter to 0.7 and the top-p value to 0.4, ensuring controlled generation diversity and quality. For each protein, we generated 100 samples and selected the pairs that exhibitd the highest structural similarity to the two target conformations. Our model successfully sampled 6/9 proteins. As shown in Figure 5, all three proteins identified by AFcluster were successfully sampled by our model, although their secondary structures remained suboptimal. The other three successful cases are presented in the Supplementary Figure S1.
@@ 4.   Concluding Remarks
Our current model may exhibit slight performance gaps in certain tasks compared to task-specific or domain-expert models, such as ESMFold2, 1. Notably, ESMFold’s success is largely attributed to its massive parameter count and extensive training dataset, which surpass even those of Llama3 8B in scale. Considering the well-known scaling laws of LLMs, we can anticipate continuous performance improvements across various tasks with the development of larger LLMs and more biological data. Beyond scaling parameters, well-established paradigms in the LLM domain, such as reinforcement learning-based alignment64 and inference-time self-improvement65, can be systematically applied to enhance ProtTeX’s performance across various protein-related tasks. Exploring these approaches will be a key focus for our future research.
@@ A.2.   Dataset
UniProt43 is a comprehensive and widely used database for protein annotation, enriched with detailed functional information and corresponding protein sequences. Many protein-related question-answering datasets focus on protein functionality. For example, Mol-Instruction19 and ProteinLMBench 42 are both constructed based on the extensive information provided by UniProt.To establish a robust protein structure database, we first construct a dataset based on UniProt. Specifically, we collect the clustered AlphaFold Protein Structure Database (AFDB) v4 dataset39, which includes 2.27 million single-chain structures predicted before July 25, 2022. Additionally, we extract protein structures from the Swiss-Prot database, filtering and curating the dataset released in May 2022, resulting in 541,327 single-chain structures. Furthermore, we incorporate experimentally determined structures from RCSB PDB41, including 551,957 single-chain structures released before October 13, 2021. In total, our dataset comprises 3.36 million protein sequences paired with structural information, forming a comprehensive resource for protein structure analysis and model training.
@@
Protein Structure Analysis Dataset(PSAD). We curate the dataset from proteinLMBench 42. The original questions required the model to generate descriptions of protein subunit composition, including the organization and interactions between subunits based on the given protein sequence. To enhance the model’s capabilities, we redesigned the prompt to enable protein structure generation following the completion of structural analysis, as illustrated in Figure 1C. This modification allows the model to achieve multimodal Chain-of-Thought (CoT) functionality.
@@ A.3.   Training Details
In our experiments, we employ both full-parameter fine-tuning and Low-Rank Adaptation (LoRA) 48 to optimize the performance of LLMs. To determine the most effective approach, we conduct ablation studies, which reveal that LoRA significantly underperforms compared to full-parameter fine-tuning in terms of task-specific metrics. Consequently, we adopt full-parameter fine-tuning as our primary training strategy. For optimization, we use the AdamW optimizer with a weight decay of 0.1. During training, we apply a cosine annealing learning rate scheduler, gradually decreasing the learning rate from 5e-6 to 1e-7, complemented by a warm-up phase with a ratio of 0.01. Training is conducted on the entire dataset with a batch size of 3 per device over 4 epochs, requiring approximately 5 days to complete on a cluster of 16 NVIDIA A100 GPUs.
@@ A.4.   Ablation details
Ablation on training strategy. We compare the LoRA (Low-Rank Adaptation) 48 training method with full-parameter fine-tuning. For full-parameter fine-tuning, we adopt the configuration previously described in Section A.3. For LoRA implementation, we configure the following parameters: LoRA target set to ‘all’, LoRA rank = 8, and alpha = 16. The learning rate is set higher than that of full-parameter fine-tuning, initialized at 1e-4 and gradually decayed to 1e-6, while all other hyperparameters remain consistent with the full-parameter fine-tuning setup. Our experimental results indicate that LoRA training exhibited significantly slower convergence in validation loss compared to full-parameter fine-tuning, as illustrated in the supplementary figures S2. Moreover, with the application of DeepSpeed Zero Stage 3 technology, LoRA does not demonstrate significant advantages in terms of memory usage or training speed across multiple GPUs compared to full-parameter fine-tuning. Consequently, we adopt full-parameter fine-tuning as our standard training paradigm in subsequent experiments.
@@
Figure S1: Multi-conformation sampling for fold-switching proteins in W. Schafer et al. 25.